# Fabian Mauthe

### v2

@@ 49 NIST - Robot Arm Position Accuracy
National Institute of Standards and Technology NIST (USA), Engineering Laboratory, Intelligent Systems Division Robotic
@@
Signals: current, position, speed Diagnosis Link
@@ 50 OpenEI - Evaluation of Building Fault
Lin Guanjing and Robin Mitchel, Lawrence Berkeley National Laboratory, (2019) Building
@@
Signals: pressure, speed, temperature Diagnosis Link
@@ 51 OpenEI - Gearbox Fault Diagnosis
Yogesh Pandy, National Renewable Energy Laboratory Mechanical component
@@
Signals: vibration Fault detection Link
@@ 52 Oxford - Oxford Battery Degradation Dataset
Signals: current, temperature, voltage Prognosis Link
@@ 53 PHM Data Challenge 2008 - Turbofan
NASA Prognostics Center of Excellence Drive technology
@@
Signals: anonymized, operating condition Prognosis Link
@@ 54 PHM Data Challenge 2009 - Gearbox Fault Detection
PHM Society, Gearbox fault detection data set, 2010 Mechanical component
@@
Signals: speed, vibration Diagnosis Link
@@ 55 PHM Data Challenge 2010 - CNC milling machine cutters
X. Li, B.S. Lim, J.H. Zhou, S. Huang, S.J. Phua, K.C. Shaw, and M.J. Er; Singapore Institute of Manufacturing Technology, 71 Nanyang Drive, Singapore 638075 School of Electrical and Electronic Engineering, Nanyang Technological University, Nanyang Avenue, Singapore 639798 Manufacturing process
@@
Signals: acoustic emission, force, vibration Prognosis Link
@@ 56 PHM Data Challenge 2011 - Anemometer Fault Detection
Creators Unknown Mechanical component
@@
Signals: direction, speed, temperature Fault detection Link
@@ 57 PHM Data Challenge 2013 - Unknown
Signals: anonymized Health assessment Link
@@ 58 PHM Data Challenge 2014 - Unknown
Signals: anonymized Health assessment Link
@@ 59 PHM Data Challenge 2015 - Plant Fault Detection
Creators Unknown Production system
@@
Signals: anonymized Diagnosis Link
@@ 60 PHM Data Challenge 2016 - Semiconductor CMP
Crystec Technology Trading GmbH Manufacturing process
@@
Signals: flow rate, operating condition, pressure, speed Prognosis Link
@@ 61 PHM Data Challenge 2017 - Bogie Vehicle
Creators Unknown Mechanical component
@@
Signals: vibration Diagnosis Link
62 PHM Data Challenge 2018 - Ion Mill in Waver Manufactoring
Kai Goebel (at this time at NASA) Production system
@@ 62 PHM Data Challenge 2018 - Ion Mill in Waver Manufactoring
Signals: current, pressure, speed Prognosis Link
@@ 63 PHM Data Challenge 2019 - Fatigue Cracks
Creators Unknown Materials
@@
Signals: vibration Prognosis Link
@@ 64 PHM Data Challenge 2020 Europe - Filtration System
Eker, O. F., Camci, F., & Jennions, I. K. (2016). Physics-based prognostic modelling of filter clogging phenomena. Mechanical Systems and Signal Processing, 75, pp. 395-412. Process technology
@@
Signals: flow rate, pressure Prognosis Link
@@ 65 PHM Data Challenge 2021 - Turbofan 2
M. Chao, C.Kulkarni, K. Goebel and O. Fink (2021). ”Aircraft Engine Runto-Failure Dataset under real flight conditions”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA (additional validation data compared to data set NASA - Turbofan engine degradation simulation data set 2 ) Drive technology
@@
Signals: anonymized, operating condition Prognosis Link
@@ 66 PHM Data Challenge 2021 Europe: SCARA-robot
Swiss Centre for Electronics and Microtechnology (CSEM) Robotic
@@
Signals: current, duration, heat slope, humidity, position, pressure, speed, temperature Diagnosis Link
@@ 67 PHM Data Challenge 2022 - Rock Drills
Epiroc Rock Drills AB, Örebro, 702 25, Sweden and Linköping University, Linköping, 581 83, Sweden Mechatronic system
@@
Signals: pressure Diagnosis Link
@@ 68 PHM Data Challenge 2022 Europe - PCB Production line
Danilo Giordano and Martino Trevisan, (2022). ”PHME Data Challenge”. European conference of the prognostics and health management society. / BITRON: Strada del Portone 95 - 10095 Grugliasco (TO), Italy, https://www.bitron.com/en Production system
@@
Signals: inspection data Diagnosis Link
@@ 69 PHM Data Challenge 2023 - Gearbox
PHM Society, Gearbox data set, 2023 Mechanical component
@@
Signals: speed, vibration Diagnosis Link
@@ 70 PHM Data Challenge 2023 Asia Pacific - Experimental Propulsion System
Japan Aerospace Exploration Agency (JAXA) Drive technology
@@
Signals: pressure Health assessment Link
@@ 71 PHM Data Challenge 2024 - Helicopter Turbine Engines
Prognostics and Health Management (PHM) Society Drive technology
@@
Signals: power, speed, temperature, torque Fault detection Link
@@ 72 PHM IEEE Data Challenge 2012 - FEMTO Bearing Data Set
FEMTO-ST Institute, Besan¸con, France Mechanical component
@@
Signals: temperature, vibration Prognosis Link
@@ 73 PHM IEEE Data Challenge 2014 - Fuel Cell
FCLAB Research Federation (FR CNRS 3539, France) Electrical component
@@
Signals: current, current density, flow rate, hygrometry, pressure, temperature, voltage Prognosis Link
@@ 74 PHM IEEE Data Challenge 2023 - Planetary Gearbox
Mälardalen University (MDU) / Mälardalen Industrial Technology Center (MITC) Mechanical component
@@
Signals: vibration Diagnosis Link
@@ 75 PHM Society - Electromechanical Ball Screw Drive
GE Research and University of Tennessee Knoxville in part by the Advanced Research Projects AgencyEnergy (ARPA-E), U.S. Department of Energy, under Award Number DEAR0001290 Mechatronic system
Electromechanical Ball Screw Drive*
Signals: current, position, speed, torque Diagnosis Link
@@ 76 ResearchGate - Driveline Unbalanced Shaft
Giacomo Barbieri and David SanchezLondono and Laura Cattaneo and Luca Fumagalli and David Romero Mechanical component
@@
Signals: vibration Diagnosis Link
@@ 77 SDOL - Diagnostics 101 bearing data
Kim, S., An, D., Choi, J-H. (2020). Diagnostics 101: A Tutorial for Fault Diagnostics of Rolling Element Bearing Using Envelope Analysis in MATLAB. Applied Sciences. 10(20):7302. doi:10.3390/app10207302 Mechanical component
@@
Signals: vibration Diagnosis Link
@@ 78 SDOL - HS Gear
Creators Unknown Mechanical component
@@
Signals: vibration Fault detection Link
@@ 79 SDOL - KAU Gear
Creators Unknown Mechanical component
@@
Signals: encoder Diagnosis Link
@@ 80 Toyota Research Institute - Battery Cycle Life
Severson, K.A., Attia, P.M., Jin, N. et al. Data-driven prediction of battery cycle life before capacity degradation. Nat Energy 4, 383–391 (2019). https://doi.org/10.1038/s41560-0190356-8 Electrical component
@@
Signals: current, temperature, voltage Prognosis Link
@@ 81 UBFC - AMPERE Detection and diagnostics of rotor and stator faults in rotating machines
Moncef Soualhi (Franche-Comté Electronique Mécanique Thermique et Optique - Sciences et Technologies (UMR 6174) (Université de Franche-Comté)), Abdenour Soualhi (Laboratoire d’Analyse des Signaux et des Processus Industriels), Thi-Phuong Khanh Nguyen, Kamal Medjaher (both Laboratoire Génie de production), Guy Clerc, Hubert Razik (both Laboratoire Ampère), doi: 10.25666/DATAUBFC-2023-03-06-03 Drive technology
@@
Signals: current, speed, vibration, voltage Diagnosis Link
@@ 82 UBFC - LASPI Detection and diagnostics of gearbox faults
Moncef Soualhi (Franche-Comté Electronique Mécanique Thermique et Optique - Sciences et Technologies (UMR 6174) (Université de Franche-Comté)), Abdenour Soualhi (Laboratoire d’Analyse des Signaux et des Processus Industriels), Thi-Phuong Khanh Nguyen, Kamal Medjaher (both Laboratoire Génie de production), Guy Clerc, Hubert Razik (both Laboratoire Ampère), doi: 10.25666/DATAUBFC-2023-03-06 . Mechanical component
@@
Signals: current, vibration, voltage Diagnosis Link
@@ 83 UBFC - METALLICADOUR Detection and diagnostics of multi-axis robot faults
Moncef Soualhi (Franche-Comté Electronique Mécanique Thermique et Optique - Sciences et Technologies (UMR 6174) (Université de Franche Comté)), Abdenour Soualhi (Laboratoire d’Analyse des Signaux et des Processus Industriels), Thi-Phuong Khanh Nguyen, Kamal Medjaher (both Laboratoire G´enie de production), Guy Clerc, Hubert Razik (both Laboratoire Ampère). Robotic
@@
Signals: current, force, position, torque, vibration Diagnosis Link
@@ 84 UCI - AI4I 2020 Predictive Maintenance
Matzka, S. (2020). “Explainable Artificial Intelligence for Predictive Maintenance Applications,” Proceedings of the Third International Conference on Artificial Intelligence for Industries (AI4I), pp. 69-74, doi: 10.1109/AI4I49448.2020.00023 . Production system
@@
Signals: speed, temperature, torque, wear Diagnosis Link
@@ 85 UCI - Condition Based Maintenance of Naval Propulsion Plants Data Set
1: DIBRIS - University of Genoa 2: School of Marine Science and Technology, Newcastle University Drive technology
@@
Signals: flow rate, injection control, pressure, speed, temperature, torque Prognosis Link

### v3

@@ 56 NIST - Robot Arm Position Accuracy
National Institute of Standards and Technology NIST (USA), Engineering Laboratory, Intelligent Systems Division
Robotic
@@
Signals: current, position, speed
Diagnosis Link
@@ 57 OpenEI - Evaluation of Building Fault
Lin Guanjing and Robin Mitchel, Lawrence Berkeley National Laboratory, (2019)
Building
@@
Signals: pressure, speed, temperature
Diagnosis Link
@@ 58 OpenEI - Gearbox Fault Diagnosis
Yogesh Pandy, National Renewable Energy Laboratory
Mechanical component
@@
Signals: vibration
Fault detection Link
59 OSF - Second-life lithium-ion battery aging dataset based on grid storage cycling
Khan, M. A., & Onori, S. (2025). Second-life lithium-ion battery aging dataset based on grid storage cycling. https://doi.org/10.17605/OSF.IO/8JNR5
Electrical component
Battery
Signals: current, impedance, voltage
Prognosis Link
@@ 60 Oxford - Oxford Battery Degradation Dataset
Signals: current, temperature, voltage
Prognosis Link
@@ 61 PHM Data Challenge 2008 - Turbofan
NASA Prognostics Center of Excellence
Drive technology
@@
Signals: anonymized, operating condition
Prognosis Link
@@ 62 PHM Data Challenge 2009 - Gearbox Fault Detection
PHM Society, Gearbox fault detection data set, 2010
Mechanical component
@@
Signals: speed, vibration
Diagnosis Link
@@ 63 PHM Data Challenge 2010 - CNC milling machine cutters
X. Li, B.S. Lim, J.H. Zhou, S. Huang, S.J. Phua, K.C. Shaw, and M.J. Er; Singapore Institute of Manufacturing Technology, 71 Nanyang Drive, Singapore 638075 School of Electrical and Electronic Engineering, Nanyang Technological University, Nanyang Avenue, Singapore 639798
Manufacturing process
@@
Signals: acoustic emission, force, vibration
Prognosis Link
@@ 64 PHM Data Challenge 2011 - Anemometer Fault Detection
Creators Unknown
Mechanical component
@@
Signals: direction, speed, temperature
Fault detection Link
@@ 65 PHM Data Challenge 2013 - Unknown
unknown
Signals: anonymized
Health assessment Link
@@ 66 PHM Data Challenge 2014 - Unknown
unknown
Signals: anonymized
Health assessment Link
@@ 67 PHM Data Challenge 2015 - Plant Fault Detection
Creators Unknown
Production system
@@
Signals: anonymized
Diagnosis Link
@@ 68 PHM Data Challenge 2016 - Semiconductor CMP
Crystec Technology Trading GmbH
Manufacturing process
@@
Signals: flow rate, operating condition, pressure, speed
Prognosis Link
@@ 69 PHM Data Challenge 2017 - Bogie Vehicle
Creators Unknown
Mechanical component
@@
Signals: vibration
Diagnosis Link
70 PHM Data Challenge 2018 - Ion Mill in Wafer Manufacturing
Kai Goebel (at this time at NASA)
Production system
@@ 70 PHM Data Challenge 2018 - Ion Mill in Wafer Manufacturing
Signals: current, pressure, speed
Prognosis Link
@@ 71 PHM Data Challenge 2019 - Fatigue Cracks
He J, Guan X, Peng T, Liu Y, Saxena A, Celaya J, Goebel K (2013). A multi-feature integration method for fatigue crack detection and crack length estimation in riveted lap joints using Lamb waves. Smart Materials and Structures. 22(10):105007.
Materials
@@
Signals: vibration
Prognosis Link
@@ 72 PHM Data Challenge 2020 Europe - Filtration System
Eker, O. F., Camci, F., & Jennions, I. K. (2016). Physics-based prognostic modelling of filter clogging phenomena. Mechanical Systems and Signal Processing, 75, pp. 395-412.
Process technology
@@
Signals: flow rate, pressure
Prognosis Link
@@ 73 PHM Data Challenge 2021 - Turbofan 2
M. Chao, C.Kulkarni, K. Goebel and O. Fink (2021). ”Aircraft Engine Run-to-Failure Dataset under real flight conditions”, NASA Ames Prognostics Data Repository, NASA Ames Research Center, Moffett Field, CA (additional validation data compared to data set NASA - Turbofan engine degradation simulation data set 2 )
Drive technology
@@
Signals: anonymized, operating condition
Prognosis Link
@@ 74 PHM Data Challenge 2021 Europe: SCARA-robot
Swiss Centre for Electronics and Microtechnology (CSEM)
Robotic
@@
Signals: current, duration, heat slope, humidity, position, pressure, speed, temperature
Diagnosis Link
@@ 75 PHM Data Challenge 2022 - Rock Drills
Epiroc Rock Drills AB, Örebro, 702 25, Sweden and Linköping University, Linköping, 581 83, Sweden
Mechatronic system
@@
Signals: pressure
Diagnosis Link
@@ 76 PHM Data Challenge 2022 Europe - PCB Production line
Danilo Giordano and Martino Trevisan, (2022). ”PHME Data Challenge”. European conference of the prognostics and health management society. / BITRON: Strada del Portone 95 - 10095 Grugliasco (TO), Italy, https://www.bitron.com/en
Production system
@@
Signals: inspection data
Diagnosis Link
@@ 77 PHM Data Challenge 2023 - Gearbox
PHM Society, Gearbox data set, 2023
Mechanical component
@@
Signals: speed, vibration
Diagnosis Link
@@ 78 PHM Data Challenge 2023 Asia Pacific - Experimental Propulsion System
Japan Aerospace Exploration Agency (JAXA)
Drive technology
@@
Signals: pressure
Health assessment Link
@@ 79 PHM Data Challenge 2024 - Helicopter Turbine Engines
Prognostics and Health Management (PHM) Society
Drive technology
@@
Signals: power, speed, temperature, torque
Fault detection Link
80 PHM Data Challenge 2025 - Aircraft Engines
Creators Unknown, data generation: NASA AGTF30 Simulation – MATLAB Executable Steady-State Solver and Linearization Tool for the AGTF30 Engine Simulation (MEXLIN-AGTF30), https://software.nasa.gov/software/LEW-20688-1. See also https://software.nasa.gov/software/LEW-19717-1. MATLAB T-MATS Toolbox – https://ntrs.nasa.gov/api/citations/20180002976/downloads/20180002976.pdf
Drive technology
Aircraft Engine*
Signals: operation conditions, pressure
Prognosis Link
81 PHM Data Challenge 2025 Asia Pacific - Predicting Cutter Flank Wear
PHM Asia Pacific
Manufacturing process
Turning machine cutter flank wear
Signals: acoustic emission, vibration
Prognosis Link
82 PHM Data Challenge 2026 Europe - Subway ticket validation door
Soualhi, M. (2026): PHME Data Challenge. European conference of the prognostics and health management society
Mechatronic system
Validation door
Signals: current, position, voltage
Prognosis Link
@@ 83 PHM IEEE Data Challenge 2012 - FEMTO Bearing Data Set
FEMTO-ST Institute, Besançon, France
Mechanical component
@@
Signals: temperature, vibration
Prognosis Link
@@ 84 PHM IEEE Data Challenge 2014 - Fuel Cell
FCLAB Research Federation (FR CNRS 3539, France)
Electrical component
@@
Signals: current, current density, flow rate, hygrometry, pressure, temperature, voltage
Prognosis Link
@@ 85 PHM IEEE Data Challenge 2023 - Planetary Gearbox
Mälardalen University (MDU) / Mälardalen Industrial Technology Center (MITC)
Mechanical component
@@
Signals: vibration
Diagnosis Link
@@ 86 PHM Society - Electromechanical Ball Screw Drive
GE Research and University of Tennessee Knoxville in part by the Advanced Research Projects Agency-Energy (ARPA-E), U.S. Department of Energy, under Award Number DEAR0001290
Mechatronic system
Ball Screw Drive* (Electromechanical)
Signals: current, position, speed, torque
Diagnosis Link
@@ 87 ResearchGate - Driveline Unbalanced Shaft
Giacomo Barbieri and David Sanchez-Londono and Laura Cattaneo and Luca Fumagalli and David Romero
Mechanical component
@@
Signals: vibration
Diagnosis Link
@@ 88 SDOL - Diagnostics 101 bearing data
Kim, S., An, D., Choi, J-H. (2020). Diagnostics 101: A Tutorial for Fault Diagnostics of Rolling Element Bearing Using Envelope Analysis in MATLAB. Applied Sciences. 10(20):7302. doi:10.3390/app10207302
Mechanical component
@@
Signals: vibration
Diagnosis Link
@@ 89 SDOL - HS Gear
Creators Unknown
Mechanical component
@@
Signals: vibration
Fault detection Link
@@ 90 SDOL - KAU Gear
Creators Unknown
Mechanical component
@@
Signals: encoder
Diagnosis Link
@@ 91 Toyota Research Institute - Battery Cycle Life
Severson, K.A., Attia, P.M., Jin, N. et al. Data-driven prediction of battery cycle life before capacity degradation. Nat Energy 4, 383–391 (2019). https://doi.org/10.1038/s41560-0190356-8
Electrical component
@@
Signals: current, temperature, voltage
Prognosis Link
@@ 92 UBFC - AMPERE Detection and diagnostics of rotor and stator faults in rotating machines
Moncef Soualhi (Franche-Comté Electronique Mécanique Thermique et Optique - Sciences et Technologies (UMR 6174) (Université de Franche-Comté)), Abdenour Soualhi (Laboratoire d’Analyse des Signaux et des Processus Industriels), Thi-Phuong Khanh Nguyen, Kamal Medjaher (both Laboratoire Génie de production), Guy Clerc, Hubert Razik (both Laboratoire Ampère), doi: 10.25666/DATAUBFC-2023-03-06-03
Drive technology
@@
Signals: current, speed, vibration, voltage
Diagnosis Link
@@ 93 UBFC - LASPI Detection and diagnostics of gearbox faults
Moncef Soualhi (Franche-Comté Electronique Mécanique Thermique et Optique - Sciences et Technologies (UMR 6174) (Université de Franche-Comté)), Abdenour Soualhi (Laboratoire d’Analyse des Signaux et des Processus Industriels), Thi-Phuong Khanh Nguyen, Kamal Medjaher (both Laboratoire Génie de production), Guy Clerc, Hubert Razik (both Laboratoire Ampère), doi: 10.25666/DATAUBFC-2023-03-06 .
Mechanical component
@@
Signals: current, vibration, voltage
Diagnosis Link
@@ 94 UBFC - METALLICADOUR Detection and diagnostics of multi-axis robot faults
Moncef Soualhi (Franche-Comté Electronique Mécanique Thermique et Optique - Sciences et Technologies (UMR 6174) (Université de Franche Comté)), Abdenour Soualhi (Laboratoire d’Analyse des Signaux et des Processus Industriels), Thi-Phuong Khanh Nguyen, Kamal Medjaher (both Laboratoire Génie de production), Guy Clerc, Hubert Razik (both Laboratoire Ampère).
Robotic
@@
Signals: current, force, position, torque, vibration
Diagnosis Link
@@ 95 UCI - AI4I 2020 Predictive Maintenance
Matzka, S. (2020). “Explainable Artificial Intelligence for Predictive Maintenance Applications,” Proceedings of the Third International Conference on Artificial Intelligence for Industries (AI4I), pp. 69-74, doi: 10.1109/AI4I49448.2020.00023 .
Production system
@@
Signals: speed, temperature, torque, wear
Diagnosis Link
@@ 96 UCI - Condition Based Maintenance of Naval Propulsion Plants Data Set
1: DIBRIS - University of Genoa 2: School of Marine Science and Technology, Newcastle University
Drive technology
@@
Signals: flow rate, injection control, pressure, speed, temperature, torque
Prognosis Link